OpenStudy (anonymous):

Write y=x-4 in standard form (Ax+By=C)

OpenStudy (therealmeeeee):

if by standard form you mean y=mx+b your equation is already in standard form gradient m=1 y intercept b=−4

OpenStudy (anonymous):

just subtract that x to the other side, and you get.... - x + y = -4 this is pretty standard, but sometimes people want A to be postitive, so you could multiply everything by -1 and get..... x - y = 4 which is the same thing, just a little less negative

OpenStudy (shamil98):

Standard form: ax +by = c move x to the other side.

OpenStudy (anonymous):

x-y=4